Comment karlsuss2 SNF 2010-11-08 15:54:24: Wec Problem

gsosa at snf.stanford.edu gsosa at snf.stanford.edu
Tue Nov 9 12:12:31 PST 2010


Intermittent with WEC head not resetting correctly. WEC not making sensor S 1.9( Z axis at bottom sensor). Checked all cables and connectors. Checked for obstructions in WEC travel. Made a 1/4 turn CCW adjustment to sensor S 1.9. Tested reset function many times with no failure. Tested basic machine operation with dummy mask and wafer-OK. Checked and calibrated exposure to 15.0 Mw/cm2.




More information about the karlsuss2-pcs mailing list